ammonia and hydrogen chloride gases react to form solid ammonium chloride. A 10.0 L reaction flask contains ammonia at 0.452 atm and 22 degree celcius, and 155 mL hydrogen chloride gas at 7.50 atm and 271 K is introduced. After the reaction occurs and temperature returns to 22 degree celcius , what is the pressure inside the flask? (Neglect the volume of the solid product)
1
Expert's answer
2020-05-07T14:16:55-0400

Use ideal gas law to determine number of moles of gases:

p*V = n*R*T, where p is pressure in Pa, V is volume in m3, n is number of moles, R = 8.314 J/(K*mole), T is temperature in K.

Initial amount of ammonia was:

n(NH3) = 0.452*101325 Pa * 0.010 m3 / 8.314 J/(K*mole) / (22+273 K) = 0.1867 mole.

Initial amount of hydrogen chloride was:

n(HCl) = 7.50*101325 Pa * 0.000155 m3 / 8.314 J/(K*mole) / 271 K = 0.0523 mole.

After the reaction:

NH3(g) + HCl(g) = NH4Cl(s)

the flask contains 0.1867 - 0.0523 = 0.1344 mole of gaseous ammonia, because ammonia was in excess. Its pressure is:

p = 0.1344 mole * 8.314 J/(K*mole) * (22+273 K) / 0.010 m3 = 32963 Pa = 0.325 atm.

Answer: final pressure is 0.325 atm